Help me please!!!
A line in the Cartesian plane passes through the points (-2,4) and 2,-1). What are the rise and the run of the line?

Answer by stanbon(48522)

A line in the Cartesian plane passes through the points (-2,4) and 2,-1). What are the rise and the run of the line?
Rise = change in Y-values = -1-4=-5
Run = corresponding change in x-values = 2-(-2)=4
Cheers,
Stan H.